# Understanding Ethical Practices in AI: Insights from a Cross-Role, Cross-Region Survey of AI Development Teams

### v1

@@
Abstract.
Recent advances in AI applications have raised growing concerns about the need for ethical guidelines and regulations to mitigate the risks posed by these technologies. In this paper, we present a mixed-method survey study - combining statistical and qualitative analyses - to examine the ethical perceptions, practices, and knowledge of individuals involved in various AI development roles. Our survey includes 414 participants from 43 countries, representing roles such as AI managers, analysts, developers, quality assurance professionals, and information security and privacy experts. The results reveal varying degrees of familiarity and experience with AI ethics principles, government initiatives, and risk mitigation strategies across roles, regions, and other demographic factors. Our findings highlight the importance of a collaborative, role-sensitive approach, involving diverse stakeholders in ethical decision-making throughout the AI development lifecycle. We advocate for developing tailored, inclusive solutions to address ethical challenges in AI development, and we propose future research directions and educational strategies to promote ethics-aware AI practices.
@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: $c$s: Software and its engineering AI Software Developmentccs: Human and Social Aspects Empirical Studies
1. Introduction
The rise of generative AI has intensified debates over its socio-economic impacts. Critics argue that these technologies can amplify issues such as discrimination, disinformation, deception, fraud, manipulation, and mistrust (Bender et al., 2021; for AI Safety, 2023a; Gebru et al., 2023a, b; The Guardian, 2023a). A central concern is the lack of accountability and transparency, as many algorithms operate as “black boxes,” obscuring how decisions are made. Additional challenges include the use of non-representative datasets and the lack of diversity within development teams, factors that may contribute to cultural or ethnic biases. Furthermore, some researchers warn that AI may undermine democratic institutions and the rule of law (New York Times, 2023; of Life Institute, 2023; Gebru et al., 2023a; The Guardian, 2023b; for AI Safety, 2023a, b; Chowdhury et al., 2021; Marcus and Reuel, 2023), calling for the establishment or reinforcement of supranational governance frameworks to address these risks (Chowdhury et al., 2021; Marcus and Reuel, 2023; Wörsdörfer, 2025; Commission, 2024).
In response to these concerns, governments worldwide have introduced regulations, initiatives, and guidelines for AI ethics, such as the EU AI Act (Commission, 2024) or the OECD AI Principles (OECD, 2024), to name a few. Despite advancements in articulating AI ethics principles (Wörsdörfer, 2025; for AI Safety, 2023a, b) and the introduction of regulatory measures aimed at mitigating AI risks (Commission, 2024; Ayyamperumal and Ge, 2024; Dong et al., 2024), significant challenges persist in developing trustworthy AI systems. These include: (a) the complexity and ambiguity of regulatory frameworks (Massey et al., 2017), (b) limited understanding of ethical principles and their practical implementation (Alomar et al., 2022; Morley et al., 2023; Sanderson et al., 2023), and (c) a lack of comprehensive guidelines and tool-supported methods to facilitate the adoption of ethical and privacy-preserving solutions (Morley et al., 2023; Jain et al., 2023; Prybylo et al., 2024).
@@ 1. Introduction
Recent research suggests that building ethical AI systems can be strengthened by fostering awareness and collaboration through human-centered AI (HAI) approaches (Lee et al., 2024; Li, 2018; Shneiderman, 2020). Several studies, through semi-structured interviews or systematic literature review, have explored how AI developers or practitioners perceive and implement AI ethics principles (Pant et al., 2024; Sanderson et al., 2023; Morley et al., 2023; Rakova et al., 2021; Vakkuri et al., 2022; Griffin et al., 2024; Holstein et al., 2019; Vakkuri et al., 2019c; Deng et al., 2022; Vakkuri et al., 2019b, a; Agbese et al., 2023). However, none of these works conducted a comprehensive user study examining how multiple AI ethics principles are understood, implemented, and collaboratively addressed across various roles within AI development teams (such as requirements analysts, product managers, developers, or testers). There is also limited analysis of how regulations, organizational contexts, and regional differences influence teams’ perceptions and ethical practices.
@@
This paper presents a large-scale mixed-methods study involving 414 participants, including members of AI development teams, researchers, and ethicists, from 43 countries. The study systematically examines their knowledge and perceptions of AI ethics and its associated principles, regulatory initiatives, and best practices for mitigating AI-related risks throughout their development activities (i.e., from AI ethics requirements specification to design, development, and testing). Participants were recruited through Prolific and online platforms such as X (formerly Twitter), Reddit AI-focused subreddits, Quora, LinkedIn, Hugging Face, and Kaggle. They represent a wide range of roles within the AI development lifecycle, including managerial positions (e.g., project and product managers, CEOs, CTOs), analysts, developers, quality assurance professionals, researchers, ethicists, and privacy and security experts. Of the 411 participants who reported their location, 201 were from North America, 118 from Europe (including 12 from outside the EU, UK, and EEA), 43 from Africa, 25 from Central and South America, 16 from Asia, four from Australia, three from the Middle East, and one from another region.
@@
Our survey comprises three sections: (1) questions about participants’ demographics and experience with AI development (e.g., education, role, organization size and type, and level of involvement with AI); (2) questions examining their perceptions and practices related to AI; and (3) questions focusing on their knowledge, experiences, and risk mitigation strategies regarding AI ethics principles, regulatory initiatives, and best practices. To our knowledge, this is the first study to provide such a comprehensive evaluation across roles, regions, and other demographic factors.
@@
– RQ2: To what extent are AI development teams familiar with AI ethics principles and governance initiatives?
– RQ3: How do practices and experiences related to AI ethics principles and governance initiatives differ across roles, geographic regions, and other demographic variables?
@@
– RQ4: What are the risk perceptions regarding AI ethics, and how do AI risk mitigation practices differ across roles, locations, and other demographic characteristics?
@@
Our analysis suggests that participants tend to associate AI with process automation and performance enhancement (RQ1). Even participants not directly involved in AI development reported integrating AI applications into their daily routines, often viewing them as beneficial for enhancing productivity and accuracy. Data protection and security concerns remain prominent, particularly among those who have not yet adopted AI in their development practices.
Furthermore, our findings reveal disparities in familiarity with AI ethics principles and initiatives across different participant demographics (RQ2). Levels of awareness vary by role, company size/type, gender, and geographic location. For example, participants in oversight roles (e.g., product managers or requirements analysts), researchers, those working in government or mid-sized companies, and female participants generally reported higher levels of familiarity with most AI ethics principles. In contrast, software testers and information security or privacy professionals were, on average, less familiar with these principles, except in areas related to data protection and the right to privacy. Similarly, participants from the EU+EEA+UK reported lower familiarity with most principles and government initiatives than their North American counterparts, with the exception of data protection, the right to privacy, and the EU AI Act. These insights underscore the need for structured ethical guidelines and the inclusion of diverse perspectives in ethical decision-making to better support the integration of ethical considerations into AI development and governance.
@@
Our results show that practices and experiences related to AI ethics principles and initiatives vary significantly across roles, organizational contexts, and demographic groups (RQ3). AI managers and requirements analysts tend to be more proactive in embedding ethical guidelines into development processes and documentation, often placing greater emphasis on user rights. Academics and researchers mainly incorporate technical aspects of AI ethics principles, such as transparency and explainability of AI, as well as fairness and justice principles. Qualitative responses also suggest that organizational size and industry sector influence ethical practices, with participants from government or mid-sized companies reporting more robust ethical frameworks than those from smaller or less structured environments. Additionally, gender differences are evident: Female participants generally place greater emphasis on ethical considerations in their development activities. These findings underscore the value of diverse perspectives and the need for cross-role communication in fostering ethical decision-making within AI development teams.
@@
Risk mitigation practices vary notably by role and demographic context (RQ4). Individuals in administrative roles, such as product managers, often focus on establishing ethical guidelines, providing AI ethics training for their employees, and communicating AI ethics principles across their teams. Along with requirements analysts, they frequently lead (ethics-related) risk and impact assessments of AI systems. Those in development roles, such as developers and testers, employ complementary strategies to mitigate biases in AI systems. For instance, developers may evaluate model outputs and incorporate diverse or representative training data, while testers concentrate on identifying issues affecting vulnerable user groups and assessing performance across diverse populations. These complementary approaches support integrating ethics principles throughout the software development lifecycle. Information security and privacy team members tend to concentrate mainly on traditional security measures, such as encryption and access control, and not on other aspects of AI ethics. Researchers rely on institutional guidelines, policies, or personal judgment when developing AI tools and conducting AI-relevant research. Yet, they report challenges in writing ethics statements due to the difficulty of addressing multiple ethical considerations and the lack of precise guidance.
2. Related Work and Background
2.1. AI Ethics Principles
In recent years, various governments and agencies have created regulations or guidelines to incorporate ethics principles in AI development. This trend began with early efforts such as the US National Science and Technology Council’s preparing for the Future of AI (Science and, U.S.) and the EU’s Digital Charter, released in 2016 (Union, 2025). Recent developments include the Biden-Harris Administration’s Blueprint for an AI Bill of Rights (Whitehouse, 2023a), the Biden White House’s Safe, Secure, and Trustworthy AI Principles (Whitehouse, 2023b), the accompanying Executive Order, and the EU’s AI Act (Commission, 2024). Aligned with these efforts, research has also proposed various AI ethics principles and guidelines (AlgorithmWatch, 2024; Wörsdörfer, 2025; Attard-Frost et al., 2022; Fjeld et al., 2020; Hagendorff, 2020; Jobin et al., 2019; Leslie et al., 2021).
@@ 2.1. AI Ethics Principles
Attard-Frost et al. (Attard-Frost et al., 2023) reviewed 47 AI guidelines and introduced the FAST principles: Fairness, Accountability, Sustainability, and Transparency. They found that fairness and accountability were mentioned in 20 of the 47 documents, sustainability in 11, and transparency in 14. A related study by Fjeld et al. (Fjeld et al., 2020) examined 36 influential AI ethics principles documents produced by various stakeholders, including governments, intergovernmental organizations, corporations, professional bodies, advocacy groups, and multi-stakeholder coalitions. Their analysis identified eight recurring themes: privacy (featured in 97% of documents), accountability (97%), safety and security (81%), transparency and explainability (94%), fairness and non-discrimination (100%), human oversight (69%), professional responsibility (78%), and the promotion of human values (69%). According to Leslie et al. (Leslie et al., 2021), core principles of AI ethics can be distilled into key values such as human dignity, individual freedom and autonomy, the prevention of harm, non-discrimination, equality and diversity, transparency and explainability in AI systems, data protection and privacy rights, accountability and responsibility, and the upholding of democracy and the rule of law.
@@
Wörsdörfer (Wörsdörfer, 2025) combines various AI ethics principles and guidelines and proposes a list of ordoliberal-inspired AI ethics principles that could form the basis for “AI guardrails” (Ayyamperumal and Ge, 2024; Dong et al., 2024) and “constitutional AI” (Abiri, 2024; Huang et al., 2024). These principles include respect for human rights, data protection and privacy, harm prevention and beneficence, non-discrimination and freedom from privileges, fairness and justice, transparency and explainability, accountability and responsibility, democracy and the rule of law, and environmental and social sustainability. A human-centered approach to AI prioritizes respect for fundamental rights and dignity, aiming to preserve human agency, control, and responsibility. Central to this philosophy is the principle of beneficence, which mandates that AI systems do no harm by avoiding discrimination, manipulation, and negative profiling while respecting personal privacy. This involves actively tackling biased algorithmic decision-making to ensure freedom from bias and protecting the most vulnerable and marginalized groups, such as children, immigrants, and ethnic minorities, thereby addressing exclusion and inequality. Adopting “explainable AI” promotes algorithmic transparency and accountability, allowing for opening black-box algorithms that enhance trust. Furthermore, integrating “ethics by design” or value-sensitive design ensures that ethical considerations such as fairness and justice are woven into the fabric of AI development. This necessitates embedding AI within participatory democracy and the rule of law, emphasizing the importance of stakeholder dialogue and public debates to navigate the complexities and challenges posed by advanced technologies. In this paper, we adopt Wörsdörfer’s (Wörsdörfer, 2025) AI ethics principles due to their comprehensive scope, which encompasses those found in other frameworks, as well as their internal coherence and practical applicability. Appendix A - Table LABEL:tab:AIEthics-Definition describes each principle in more detail.
2.2. Ethics in AI Development
Development teams’ understanding and challenges in implementing ethics principles in AI systems have been investigated through user studies with developers and practitioners (Pant et al., 2024; Sanderson et al., 2023; Morley et al., 2023; Rakova et al., 2021; Vakkuri et al., 2022; Griffin et al., 2024; Holstein et al., 2019; Vakkuri et al., 2019c; Deng et al., 2022; Vakkuri et al., 2019b, a; Agbese et al., 2023). Pant et al. (Pant et al., 2024) conducted a grounded theory analysis of 38 academic papers regarding practitioners’ views on AI ethics; they emphasized that cultural and personal biases among practitioners can unknowingly influence AI outcomes, underscoring the need for more diverse teams, more explicit role definitions, and structured approaches to self-reflection and ethical decision-making. They found that many AI practitioners are uncertain about who holds responsibility for moral transgressions, leading developers to shift accountability to users or upper management. They also observe limited awareness of some AI ethics principles like fairness, accountability, and privacy, but not the gaps between them and their implementation in practice.
These findings are supported by Sanderson et al. (Sanderson et al., 2023) and Morley et al. (Morley et al., 2023), who found developers are not very familiar with principles related to human-centered values, fairness, and accountability. In interviews and surveys with developers about their familiarity, opinions, and challenges in implementing AI ethics principles, they cited trade-offs between utility and explainable AI systems, resource limitations, lack of practical tools and guidelines, and insufficient organizational support. Rakova et al. (Rakova et al., 2021) found that developers face organizational obstacles, such as misaligned incentives and a lack of accountability, which prevent the proactive integration of ethics principles. Vakkuri et al. (Vakkuri et al., 2022) uncovered that many companies struggle with integrating AI ethics into existing workflows, particularly in balancing ethics principles with efficiency and business goals, leading to the deprioritization of moral considerations. Griffin et al. (Griffin et al., 2024) and Agbese et al. (Agbese et al., 2023) complemented these findings through interviews with AI developers, showing that while most are aware of the existing issues, they lack formal training and support to navigate them consistently and often rely on informal advice from colleagues due to insufficient organizational frameworks.

### v2

@@
Abstract
Recent advances in AI applications have raised growing concerns about the need for ethical guidelines and regulations to mitigate the risks posed by these technologies. In this paper, we present a mixed-methods survey study - combining statistical and qualitative analyses - to examine the ethical perceptions, practices, and knowledge of individuals involved in various AI development roles. Our survey comprises 414 participants from 43 countries, representing various roles such as AI managers, analysts, developers, quality assurance professionals, and information security and privacy experts. The results reveal varying degrees of familiarity and experience with AI ethics principles, government initiatives, and risk mitigation strategies across roles, regions, and other demographic factors. Our findings underscore the importance of a collaborative, role-sensitive approach that involves diverse stakeholders in ethical decision-making throughout the AI development lifecycle. We advocate for developing tailored, inclusive solutions to address ethical challenges in AI development, and we propose future research directions and educational strategies to promote ethics-aware AI practices.
keywords:
@@ Abstract
1 Introduction
The rise of generative AI has intensified debates over its socio-economic impacts. Critics argue that these technologies can amplify issues such as discrimination, disinformation, deception, fraud, manipulation, and mistrust [bender2021dangers, center2023examples, gebru2023statement, gebru2023considerations, guardian2023drake]. A central concern is the lack of accountability and transparency, as many algorithms operate as “black boxes,” obscuring how decisions are made. Additional challenges stem from the use of non-representative datasets and insufficient diversity in development teams, both of which can contribute to cultural or ethnic biases. Some researchers further warn that AI may undermine democratic institutions and the rule of law [nyt2023godfather, future2023pause, gebru2023statement, guardian2023laws, center2023examples, center2023statement, chowdhury2021developers, marcus2023world], calling for the establishment or reinforcement of (supra-)national governance frameworks to address these risks [chowdhury2021developers, marcus2023world, worsdorfer2023aibill, AIAct2024]. In response, governments worldwide have introduced regulations, initiatives, and ethical guidelines, such as the EU AI Act [AIAct2024], in an effort to mitigate emerging harms. Yet, despite progress in articulating AI ethics principles [worsdorfer2023aibill, center2023examples, center2023statement] and implementing regulatory measures aimed at reducing AI risks [AIAct2024, ayyamperumal2024current, dong2024building], substantial obstacles remain in developing trustworthy AI systems. These include: (a) the complexity and ambiguity of existing regulatory frameworks [MHG17], (b) a persistent gap between ethical principles and real-world practice, worsened by role ambiguity and limited understanding of how to operationalize ethics [pant2024ethics, griffin2024ethical, morley2023operationalising, sanderson2023ai, Alomar2022, vakkuri2020aiethics, vakkuri2022how, khan2023aiethics, kawakami2024studyingup], and (c) the absence of comprehensive guidelines and tool-supported methods that facilitate the adoption of ethical and privacy-preserving solutions [morley2023operationalising, jain2023, prybylo2024evaluating, khan2023aiethics].
@@ 1 Introduction
Recent research suggests that fostering collaboration through human-centered AI (HAI) approaches [hao2024idontknow, li2018ai, shneiderman2020bridging] can strengthen the development of AI ethics systems. While several studies, drawing on semi-structured interviews or systematic literature review, have examined how AI developers or practitioners perceive and implement AI ethics principles [pant2024ethics, sanderson2023ai, morley2023operationalising, rakova2021where, vakkuri2022how, griffin2024ethical, holstein2019improving, vakkuri2019ethically, deng2022understanding, vakkuri2019implementing, vakkuri2019ethicallyaligned, agbese2023implementing], gaps remain. Notably, no prior work has conducted a comprehensive user study that investigates how multiple AI ethics principles are understood, implemented, and collaboratively addressed across the diverse roles within AI development teams (e.g., requirements analysts, product managers, developers, or testers), or how these roles/factors influence familiarity with and attention to ethical considerations. Moreover, little is known about how regulations, organizational contexts, and regional differences shape teams’ perceptions and ethical practices.
@@
This paper presents a large-scale mixed-methods study involving 414 participants, from 43 countries. The study systematically examines their knowledge and perceptions of AI ethics and its associated principles, regulatory initiatives, and best practices for mitigating AI-related risks throughout their development activities (i.e., from AI ethics requirements specification to design, development, and testing). Participants were recruited through Prolific and online platforms such as X (formerly Twitter), Reddit AI-focused subreddits, Quora, LinkedIn, Hugging Face, and Kaggle. They represent a wide range of roles within the AI development lifecycle, including managerial positions (e.g., project and product managers, CEOs, CTOs), analysts, developers, quality assurance professionals, researchers, ethicists, and privacy and security experts. Of the 411 participants who reported their location, 201 were from North America, 118 from Europe (including 12 from outside the EU, UK, and EEA), 43 from Africa, 25 from Central and South America, 16 from Asia, four from Australia, three from the Middle East, and one from another region.
@@
Our survey comprises three sections: (1) questions about participants’ demographics and experience with AI development (e.g., education, role, organization size and type, and level of involvement with AI); (2) questions examining their perceptions and practices related to AI; and (3) questions focusing on their knowledge, experiences, and risk mitigation strategies regarding AI ethics principles, regulatory initiatives, and best practices. To the best of our knowledge, this is the first study to provide such a comprehensive evaluation across roles, regions, and other demographic factors.
@@
– RQ2: To what extent are AI development teams familiar with AI ethics principles and governance initiatives, and what are the predictive factors for such familiarity?
@@
– RQ3: How do practices and perceptions regarding AI ethics principles and initiatives differ across roles, and other demographics, and what are the predictive factors?
– RQ4: What are the risk perceptions regarding AI ethics, and how do AI risk mitigation practices differ across roles and other demographic characteristics?
@@
Our analysis suggests that participants tend to associate AI with process automation and performance enhancement (RQ1). Even those not involved in AI development use AI in their daily routines, often viewing it as beneficial for enhancing productivity and accuracy. Data protection and security concerns remain prominent, particularly among those who have not yet adopted AI in their development practices.
Our findings reveal disparities in familiarity with AI ethics principles and governance initiatives across different demographics (RQ2). Levels of awareness vary by role, company size/type, gender, and location. For example, participants in oversight roles (e.g., product managers or requirements analysts), researchers, those working in government or mid-sized companies, and female participants generally reported higher levels of familiarity with most AI ethics principles. Although there are notable differences between roles, our Intra-Class Correlation (ICC) revealed a higher variability within roles regarding familiarity with ethics principles than between them. The Least Absolute Shrinkage and Selection Operator (LASSO) analysis [tibshirani1996lasso] showed that familiarity with AI governance initiatives, particularly the EU AI Act, is the predictive factor most strongly associated with a higher familiarity with AI ethics principles.
@@
Our results show that practices and experiences related to AI ethics principles and initiatives vary significantly across roles, organizational contexts, and demographic groups (RQ3). AI managers and requirements analysts tend to be more proactive in embedding ethics guidelines into development processes and documentation, often placing greater emphasis on user rights. Academics and researchers primarily incorporate the technical aspects of AI ethics principles, like transparency and explainability of AI, and fairness and justice. Role is a major predictor when considering AI ethics principles in practice, according to our LASSO and mixed-effects model analysis. Qualitative responses also suggest that organizational size and industry sector influence ethical practices, with participants from government or mid-sized companies reporting more robust ethical frameworks than those from smaller or less structured environments. Additionally, gender differences are evident: Female participants generally place greater emphasis on ethical considerations in their development activities.
@@
Risk mitigation practices vary notably by role and demographic context (RQ4). Individuals in administrative roles, such as product managers, often focus on establishing clear ethical guidelines, providing AI ethics training to their employees, and communicating AI ethics principles across their teams. Along with requirements analysts, they frequently lead (ethics-related) risk and impact assessments of AI systems. Those in development roles employ complementary strategies, such as evaluating model outputs or assessing model performance across diverse populations, to mitigate biases in AI systems. Information security and privacy team members tend to focus mainly on traditional security measures, such as encryption and access control, rather than other aspects of AI ethics. Researchers rely on institutional guidelines, policies, or personal judgment when developing AI tools and conducting AI-relevant research. Yet, they report challenges in writing ethics statements due to the difficulty of addressing multiple ethical considerations and the lack of precise guidance.
Overall, these findings underscore the importance of structured ethics guidelines and robust cross-role communication to ensure that diverse perspectives inform ethical decision-making and support the effective integration of ethics into AI development. We provide detailed research and educational directions based on these findings.
2 Related Work and Background
2.1 AI Ethics Principles
In recent years, various governments and agencies created regulations or guidelines to incorporate ethics principles in AI development [NSTC2017, EUDigital, Biden-Harris-1, Biden-Harris-2], including the EU’s AI Act [AIAct2024]. Research also proposed various AI ethics principles and guidelines [algorithmwatch2024global, worsdorfer2023aibill, attard2022ethics, fjeld2020principled, hagendorff2020ethics, jobin2019global, leslie2021artificial].
@@ 2.1 AI Ethics Principles
Attard-Frost et al. [attard2023ethics] reviewed 47 AI guidelines and introduced the FAST principles: Fairness, Accountability, Sustainability, and Transparency. They found that fairness and accountability were mentioned in 20 of the 47 documents, sustainability in 11, and transparency in 14. Analyzing 36 influential AI ethics principles documents produced by various stakeholders, such as governments, corporations, professional bodies, etc reveals eight recurring themes [fjeld2020principled]: privacy (featured in 97% of documents), accountability (97%), safety and security (81%), transparency and explainability (94%), fairness and non-discrimination (100%), human oversight (69%), professional responsibility (78%), and the promotion of human values (69%). Leslie et al. [leslie2021artificial] categorize core AI ethics principles into key values like human dignity, individual freedom and autonomy, the prevention of harm, non-discrimination, equality and diversity, transparency and explainability in AI systems, data protection and privacy rights, accountability and responsibility, and the upholding of democracy and the rule of law.
@@
Wörsdörfer [worsdorfer2023aibill] combines various AI ethics principles and guidelines and proposes a list of ordoliberal-inspired AI ethics principles that could form the basis for “AI guardrails” [ayyamperumal2024current, dong2024building] and “constitutional AI” [abiri2024public, huang2024collective]. These principles include respect for human rights, data protection and privacy, harm prevention and beneficence, non-discrimination and freedom from privileges, fairness and justice, transparency and explainability, accountability and responsibility, democracy and the rule of law, and environmental and social sustainability. A human-centered approach to AI prioritizes respect for fundamental rights and dignity, aiming to preserve human agency, control, and responsibility. Central to this philosophy is the principle of beneficence, which mandates that AI systems do no harm by avoiding discrimination, manipulation, and negative profiling while respecting personal privacy. This involves actively tackling biased algorithmic decision-making to ensure freedom from bias and protecting the most vulnerable and marginalized groups, thereby addressing exclusion and inequality. Adopting “explainable AI” promotes algorithmic transparency and accountability, allowing for opening black-box algorithms that enhance trust. Integrating “ethics by design” ensures that ethical considerations such as fairness and justice are woven into the fabric of AI development. This necessitates embedding AI within participatory democracy and the rule of law, emphasizing the importance of stakeholder dialogue and public debates to navigate the complexities posed by AI technologies.
In this paper, we adopt Wörsdörfer’s [worsdorfer2023aibill] AI ethics principles due to their comprehensive scope, which encompasses those found in other frameworks, as well as their practical applicability (see Appendix B - Table LABEL:tab:AIEthics-Definition for more detail).
2.2 Ethics in AI Development
Several studies examined developers’ understanding and challenges, gaps in the practices, and organizational structures in implementing ethics principles in AI systems through user studies with developers and practitioners [sanderson2023ai, vakkuri2022how, vakkuri2019implementing, khan2023aiethics, reuel2025responsibleai, vakkuri2020aiethics, pant2024ethics, morley2023operationalising, rakova2021where, griffin2024ethical, holstein2019improving, vakkuri2019ethically, deng2022understanding, vakkuri2019ethicallyaligned, agbese2023implementing, pant2024ethicsai].
Vakkuri et al. [vakkuri2020aiethics, vakkuri2022how] investigated developers’ awareness and implementation of AI ethics principles. They found that developers often treat AI features like other software functionalities and that existing AI ethics guidelines have a limited impact on their practices. They further observed that many companies struggle to integrate AI ethics into existing workflows, particularly in balancing ethics principles with efficiency and business goals, leading to the deprioritization of moral considerations. Pant et al. [pant2024ethicsai, pant2024ethics] examined AI developers’ awareness of AI ethics principles through a user study with 100 participants and analysis of 38 academic papers. They identified that awareness varied demographically, as those with more experience, a Ph.D., or who identified as female reported higher awareness, and it was mainly influenced by workplace policies, personal interest, and media coverage. The study revealed that cultural and personal biases shape AI outcomes, while unclear role boundaries contribute to uncertainty over responsibility, often shifting accountability to users or upper management. Participants were most familiar with data protection and privacy, but had a limited understanding of human-centered principles like fairness, accountability, transparency, and explainability, which aligns with the findings of Sanderson et al. [sanderson2023ai], and Morley et al. [morley2023operationalising]. Human-centered values were viewed as the hardest to implement or uphold due to inconsistent perceptions, limited knowledge and resources, a lack of practical tools and guidelines, insufficient organizational support, inherent human biases, and trade-offs between utility and explainable AI systems; similar to [rakova2021where]. The study emphasized the importance of diverse teams, explicit role definitions, and structured approaches to ethical decision-making. Griffin et al. [griffin2024ethical], and Agbese et al. [agbese2023implementing] complemented these findings through interviews with AI developers, showing that while most are aware of the existing issues, they lack formal training and support to navigate them consistently and often rely on informal advice from colleagues due to insufficient organizational guidelines.